Given a rectangle that is 36 cm longer than it is wide. Express the perimeter (in cm) of a rectangle in terms of its width .
Answer by mananth(16946)

let width be w
length = w+36
Perimeter P = 2*(L+W)
P=2(w+2+36)
P=2(2w+36)
P=2*2(w+18)
P=4(w+18)